The LMNO partnership has the following balance sheet at January 1, 2020, prior to the admissions of new partner , Paul.

Larry, Capital $52,000

Mike, Capital 104,000

Nancy, Capital 234,000

Oliver, Capital 312,000

Required:

A) Paul contributes $98,000 into the partnership for a 25% interest. The four original partners share profits and losses equally. Using the bonus method, determine the balances for each of the five partners after Paul joins the partnership.

B) Paul contributed $284,000 into the partnership for a 20% interest. Goodwill was to be recorded . The four original partners shared profits and losses equally. After Paul made his investment, what were the individual capital balances?
Answer #1

Requirement A

Total capital of firm after admission = 52000+104000+234000+312000+98000 = $800,000

Paul's interest = 800000×25% = $200,000

Shortfall in Paul's contribution = 200000-98000 = $102,000 to be borne by remaining partners equally, i.e., 102000÷4 = $25,500 each.

Particulars Larry Mike Nancy Oliver Paul
Beginning Balance $52,000 $104,000 $234,000 $312,000
Cash bought in by Paul $98,000
Shortfall in Paul contribution ($25,500) ($25,500) ($25,500) ($25,500) $102,000
Ending balance $26,500 $78,500 $208,500 $286,500 $200,000

Requirement B

Total capital of firm before admission = 52000+104000+234000+312000 = $702,000

The above capital accounts for 80% after Paul's admission.

Hence, total capital of the firm after admission = 702,000÷80% = $877,500

Total capital contributions after admission = 52000+104000+234000+312000+284000 = $986,000

Excess Contribution/ Goodwill = 986,000-877500 = $108,500

This Goodwill of $108,500 to be enjoyed by remaining partners equally, i.e., 108500÷4 = $27,125 each.

Particulars Larry Mike Nancy Oliver Paul
Beginning Balance $52,000 $104,000 $234,000 $312,000
Cash bought in by Paul $284,000
Shortfall in Paul contribution $27,125 $27,125 $27,125 $27,125 ($108,500)
Ending balance $79,125 $131,125 $261,125 $339,125 $177,500